OpenStudy (anonymous):

what is the domain? y = √3x + 3 x ≤ – 1 x > 1 x ≥ –1

OpenStudy (jlastino):

Can there be a square root of a negative number?

OpenStudy (anonymous):

\[\sqrt{3x+3}\] so \[3x+3\geq 0\] \[3x\geq -3\] \[x\geq -1\]

OpenStudy (anonymous):

\[\sqrt{-1}=i\]

OpenStudy (anonymous):

\[\sqrt{-3}=3i\]